# Contextual Stochastic Optimization for Order Fulfillment Optimization

### v1

@@ 4.4.1 Contextual Features Selection
    Order-level features: This category includes customer location, SKU weight and dimensions, SKU type, quantities, release hour, release day, and planned lead time (i.e., the difference between the promised delivery date and the order release time). These factors can influence deviations through logistical complexity, handling requirements, and time-sensitive elements affecting the order’s processing and delivery schedule.
@@
    Fulfillment center features: This includes coordinates, location type, on-hand capacity, and on-hand inventory. The geographical location and type of fulfillment center can affect the efficiency of order processing, while capacity and inventory levels determine the speed and reliability of fulfilling orders.
@@
    Carrier-related features: This includes the level of service, promised time in transit, shipping charges, and carrier zone. These factors impact the carrier ability to meet delivery deadlines, with service level and transit times directly affecting deviations, while shipping charges and zones reflect logistical challenges and potential bottlenecks.
4.4.2 Probabilistic Multi-Class Classification
Since deviations measured in days only take on discrete values in the dataset, they inherently follow a discrete probability distribution. Thus, the random variable $\tilde{d}$ has a finite discrete support. Let $C$ be the total number of possible values $\tilde{d}$ can take, and let $\xi_{c}$ be the $c$-th possible value for $c\in[C]$. By definition, $\tilde{d}\in\{\xi_{1},\xi_{2},\ldots,\xi_{C}\}$. This discrete nature allows the conditional distribution learning problem to be framed as a multi-class classification (MC-CLF) problem. Under MC-CLF, each class represents a possible delivery deviation value in number of days, and the output class probabilities correspond directly to the probability mass function of the discrete distribution. For a new sample with covariates vector $\mathbf{s}$, let the conditional probability of class $c$ be $p_{c}(\mathbf{s})=\mathbb{P}(d=\xi_{c}\mid\mathbf{s})$, with the corresponding predicted probability from MC-CLF denoted as $\hat{p}_{c}(\mathbf{s})$.
Multinomial logistic regression (MLR) and classification trees-based models are considered to solve MC-CLF. Specifically, MLR is a classic classifier that aims to estimate the class probabilities by maximizing the log-likelihood function, which is the negative of the log loss function. The log loss is defined as $-\sum_{m=1}^{M}\sum_{c=1}^{C}\mathbf{1}\{d_{m}=\xi_{c}\}\log(\hat{p}_{c}(\mathbf{s}_{m})).$ On the other hand, classification trees use a nonparametric approach that recursively partitions the data into subsets based on feature values in a top-down manner, resulting in a tree-like model of decisions.
Let $\mathcal{R}_{1},\ldots,\mathcal{R}_{R}$ be the regions partitioned by a classification tree. Given covariates $\mathbf{s}$, the predicted probability for class $c$ in a classification tree can be expressed as:
@@ 4.4.2 Probabilistic Multi-Class Classification
Moreover, while the proposed classification models excel at predicting class labels, their predicted probabilities may not be well calibrated (Niculescu-Mizil and Caruana 2005). Therefore, probability calibration is necessary to obtain accurate conditional distributional forecasts. To address this, a cross-validation approach is adopted to fit an isotonic regression calibrator (Zadrozny and Elkan 2002). In each split, the classification model is first trained on the training set, and then its predictions on the calibration set are used to fit an isotonic regression model. Let $\hat{\tilde{p}}(\cdot)$ be the vector of calibrated probabilities and $\hat{p}(\cdot)$ be the vector of uncalibrated probabilities for each class. Isotonic regression takes the vector of uncalibrated probabilities as input and solves the following problem to determine a vector of monotonically increasing calibrated probabilities:
$\min_{\hat{\tilde{p}}(\cdot)}$ $\sum_{m=1}^{M}(d_{m}-\hat{\tilde{p}}(\mathbf{s}_{m}))^{2}$ (20)
s.t. $\hat{\tilde{p}}(\mathbf{s}_{m})\geq\hat{\tilde{p}}(\mathbf{s}_{m^{\prime}})\quad\forall m^{\prime}\neq m:\hat{p}(\mathbf{s}_{m})\geq\hat{p}(\mathbf{s}_{m^{\prime}})$
@@
To measure the accuracy of probabilistic predictions for MC-CLF model, the paper adopts the log loss and the Brier score. The latter is defined as the average squared error between predicted probabilities and actual outcomes $\sum_{m=1}^{M}\sum_{c=1}^{C}(\mathbf{1}\{d_{m}=\xi_{c}\}\hat{p}_{c}(\mathbf{s}_{m}))^{2}$.
For both metrics, lower values indicate better probabilistic prediction performance.
4.4.3 Quantile Regression
Unlike multi-class classification, quantile regression (QR) can directly estimate the cumulative distribution function (CDF) of the conditional distribution without assuming a discrete probability distribution. The CDF of the uncertain deviation $\tilde{d}$ is given by $F(u\mid\mathbf{s})=\mathbb{P}(d\leq u\mid\mathbf{s})$. Standard QR models aim at minimizing the pinball loss function, estimating different quantiles independently. To approximate the conditional distribution, these models typically calculate a series of conditional quantiles, forming a discretized approximation of the CDF. However, this approach can lead to the quantile crossing issue, where predicted quantiles do not maintain their proper order, resulting in an invalid distribution. Various methods exist to avoid quantile crossing, but tree-based quantile regression models have been shown to outperform standard QR models in distributional forecasting, particularly in delivery time performance settings (Salari et al. 2022). These tree-based methods do not suffer from the issue of quantile crossing. Therefore, this paper focuses on a regression tree-based model and a quantile random forest (QRF) model which leverages an ensemble of regression trees.

### v2

@@ 4.3 Contextual Sample Average Approximation (C-SAA)
    Candidate generation.
    1. (a)
      Draw $N_{1}$ independent cost vector samples $\{\bm{\xi}^{n}\}_{n=1}^{N_{1}}\sim\mathcal{M}(\mathbf{s})$, where $\bm{\xi}^{n}=(\xi^{n}_{iv})_{i\in\mathcal{I},\;v\in\mathcal{V}}$.
    2. (b)
      Solve
      $\mathbf{z}_{i}\;=\;\arg\min_{\mathbf{z}\in\mathcal{Z}}\;\frac{1}{N_{1}}\sum_{n=1}^{N_{1}}g(\mathbf{z},\bm{\xi}^{n}),\quad i=1,\dots,Q,$
      independently $Q$ times to obtain candidate solutions $\{\mathbf{z}_{1},\dots,\mathbf{z}_{Q}\}$ with sample‐average objectives $\{f_{1},\dots,f_{Q}\}$.
    3. (c)
      Compute the average objective $\underaccent{\bar}{f}=\tfrac{1}{Q}\sum_{i=1}^{Q}f_{i}$, which is a statistical lower bound on the true optimal value $f^{*}$ because $\underaccent{\bar}{f}$ is an unbiased estimator of $\mathbb{E}_{{\tilde{\mathbf{c}}\sim\mathbb{P}(\mathbf{c}\mid\mathbf{s})}}[\underaccent{\bar}{f}]$ and $\mathbb{E}_{{\tilde{\mathbf{c}}\sim\mathbb{P}(\mathbf{c}\mid\mathbf{s})}}[\underaccent{\bar}{f}]\leq f^{*}$.
@@
    Candidate evaluation.
    1. (a)
      For each $\mathbf{z}_{i}$, draw a larger set of evaluation samples $\{\bm{\xi}^{n}\}_{n=1}^{N_{2}}\sim\mathcal{M}(\mathbf{s})$, with $N_{2}\gg N_{1}$.
    2. (b)
      Compute the re‐estimated objectives $\bar{f}(\mathbf{z}_{i})=\tfrac{1}{N_{2}}\sum_{n=1}^{N_{2}}g(\mathbf{z}_{i},\bm{\xi}^{n})$.
    3. (c)
      Each $\bar{f}(\mathbf{z}_{i})$ is an upper bound on $f^{*}$ since each candidate solution is feasible; the difference $\bar{f}(\mathbf{z}_{i})-\underaccent{\bar}{f}$ approximates the optimality gap.
    4. (d)
      Select the candidate $\mathbf{z}^{*}$ with the smallest estimated gap.
Discussion. Unlike the reweighting approach of Bertsimas and Kallus (2020), which leverages complete historical samples for each fully-observed cost vector, C–SAA relies solely on observational data (i.e., only the costs of the selected options are available). Attempting to form all possible cost vector combinations from partial observations would lead to a combinatorial explosion as $|\mathcal{V}|$ and $|\mathbb{D}|$ grow, making such method intractable in large, high‐dimensional selection spaces.
4.4 Contextual Robust Optimization (C-RO)
The performance of C-SAA largely hinges on the accuracy of the estimated contextual distribution. Robust optimization provides a hedge against misspecification of the contextual distribution by guarding against worst–case scenarios, making it a viable approach for contextual stochastic optimization. Let $\mathcal{U}(\mathbf{s})\subset\mathbb{R}^{|\mathcal{I}|\times|\mathcal{V}|}$ represent a contextual uncertainty set for the cost vector $\tilde{\mathbf{c}}$. The Contextual Robust Optimization (C-RO) problem is
$\min_{\mathbf{z}\in\mathcal{Z}}\max_{\bm{\xi}\in\mathcal{U}(\mathbf{s})}g(\mathbf{z},\bm{\xi}).$ (11)
Two constructions of $\mathcal{U}(\mathbf{s})$ are considered: a discrete sampling–based set $\mathcal{U}_{d}(\mathbf{s})$ and a budgeted interval set $\mathcal{U}_{b}(\mathbf{s},B)$.
4.4.1 Contextual Discrete Uncertainty Set
Problem (11) can be reformulated into a single-level problem by enumerating every possible realization of the random vector $\tilde{\mathbf{c}}$, a standard technique in robust optimization. In principle, a discrete distribution can be generated as mentioned before, by combining all combinations of the observed data points. However, this will make the problem intractable as discussed.
To retain tractability, a sampling-based discrete uncertainty set in the same spirit of C-SAA is introduced. Let $\mathcal{U}_{d}(\mathbf{s})=\{\bm{\xi}^{1},\ldots,\bm{\xi}^{N}\}$ denote a collection of $N$ deviation scenarios drawn from $\mathcal{M}(\mathbf{s})$. The robust counterpart of (11) becomes
[…]
s.t. $t\geq g(\mathbf{z},\bm{\xi}^{n}),\quad n=1,\dots,N.$
4.4.2 Contextual Budgeted Interval Uncertainty Set
An alternative uncertainty set is the budgeted interval set $\mathcal{U}_{b}(\mathbf{s},B)$ (Bertsimas and Sim 2004), which is defined by a robust budget $B$. This budget limits the extent to which limits how many components of the uncertain parameter $\tilde{\mathbf{c}}$ can simultaneously deviate to their upper bounds.
Rather than estimating the full joint distribution, only the conditional lower and upper quantiles at levels $q_{\rm low}$ and $q_{\rm up}$ are needed to define a $(q_{\rm low},q_{\rm up})$ prediction interval for each option.
For each option $v\in\mathcal{V}$, let $\underline{\bm{\xi}}_{v}=\bigl{(}\underline{\xi}_{iv}\bigr{)}_{i\in\mathcal{I}}$ and $\overline{\bm{\xi}}_{v}=\bigl{(}\overline{\xi}_{iv}\bigr{)}_{i\in\mathcal{I}}$ denote the estimated conditional $q_{\rm low}$- and $q_{\rm up}$-quantiles extracted from $\mathcal{M}_{v}(\mathbf{s}_{v})$:
$\underline{\bm{\xi}}_{v}=\inf\{\mathbf{u}:\mathbb{P}_{v}(\mathbf{c}_{v}\leq\mathbf{u}\mid\mathbf{s}_{v})\geq q_{\rm low}\},\quad\overline{\bm{\xi}}_{v}=\inf\{\mathbf{u}:\mathbb{P}_{v}(\mathbf{c}_{v}\leq\mathbf{u}\mid\mathbf{s}_{v})\geq q_{\rm up}\}.$
Given a budget $B$, the budgeted interval uncertainty set is
$\mathcal{U}_{b}(\mathbf{s},B)=\Bigl{\{}\bm{\xi}:\;\xi_{iv}=\underline{\xi}_{iv}+\delta_{iv}\,(\overline{\xi}_{iv}-\underline{\xi}_{iv}),\;\sum_{i,v}\delta_{iv}\leq B,\;\delta_{iv}\in[0,1]\Bigr{\}}.$
With a fixed $\mathbf{z}$, the inner maximization problem in (11) is
$\max_{\delta_{iv}}\quad$ $\sum_{i,v}(\overline{\xi}_{iv}-\underline{\xi}_{iv})z_{iv}\delta_{iv}$ (13)
s.t. $\sum_{i,v}\delta_{iv}\leq B$ (14)
$\delta_{iv}\in[0,1],\quad\forall i,v.$ (15)
Let $y_{j}\in\{0,1\}$ and $w_{iv}^{j}\geq 0$ be auxiliary variables. Denote by $\pi\geq 0$ and $\lambda_{iv}\geq 0$ the dual variables corresponding to Constraints (14) and (15), respectively. Finally, let $M$ be a sufficiently large constant. A valid choice for the “big‑M” constant is $M\;=\;\max_{i,v,j}\bigl{\{}\,b_{iv}^{n}-b_{iv}^{1},\;b_{iv}^{\,n-1},\;(\underline{\xi}_{iv}+a_{iv}^{j})\,b_{iv}^{n}\bigr{\}}.$ By taking the dual of the inner maximization problem and introducing $y_{j}$ and $w_{iv}^{j}$, a robust counterpart of (11) is obtained as
[…]
s.t. $\pi+\lambda_{iv}\geq(\overline{\xi}_{iv}-\underline{\xi}_{iv})z_{iv},\quad\forall i,v$
$z_{iv}\leq b^{j}_{iv}+M(1-y_{j}),\quad\forall i,v,j$
$z_{iv}\geq b^{j-1}_{iv}-M(1-y_{j}),\quad\forall i,v,j$
$w_{iv}^{j}\geq(\underline{\xi}_{iv}+a_{iv}^{j})z_{iv}-M(1-y_{j}),\quad\forall i,v,j$
$y_{j}\in\{0,1\},~{}w_{iv}^{j}\geq 0,~{}\pi\geq 0,~{}\lambda_{iv}\geq 0.$
4.5 Performance Guarantees on C-SAA & C-RO
C–SAA replaces the unknown cost distribution with Monte-Carlo samples generated by a learned contextual oracle and minimizes the resulting empirical objective. Under standard regularity conditions, the SAA objective is statistically consistent: its optimal value converges exponentially fast to the true optimum as $N_{1}$ grows (Kleywegt et al. 2002, Shapiro et al. 2021). When the oracle is well-calibrated and $N_{1}$ is large, C–SAA thus attains the smallest expected cost. However, it remains sensitive to distributional misspecification and finite-sample error, which can expose the solution to rare but severe tail events.
C–RO mitigates this risk by optimizing the worst-case cost over a prescribed uncertainty set. Let a feasible solution $\mathbf{z}$ satisfy $\max_{\bm{\xi}\in\mathcal{U}(\mathbf{s})}g(\mathbf{z},\bm{\xi})\leq t.$ For a discrete uncertainty set built from $N$ samples (C-RO-D), one obtains with probability $1-\delta$ that
$\Pr\bigl{(}g(\mathbf{z},\tilde{\mathbf{c}})>t\bigr{)}\;\leq\;\epsilon$
provided
[…]
samples are drawn (Calafiore and Campi 2006, Campi and Garatti 2008, Bertsimas et al. 2021). Although this bound grows linearly in the dimension of the uncertainty $|\mathcal{V}||\mathcal{I}|$, it too can degrade under distributional misspecification. Under a budgeted-interval uncertainty set (C-RO-B), the violation probability instead satisfies
$\Pr\bigl{(}g(\mathbf{z},\tilde{\mathbf{c}})>t\bigr{)}\;\leq\;\exp\!\Bigl{(}-\frac{B^{2}}{2\,|\mathcal{V}||\mathcal{I}|}\Bigr{)},$
so that the budget parameter $B$ directly controls the trade-off between average-case performance and robustness (Bertsimas and Sim 2004). This tunable flexibility makes C–RO-B an attractive middle ground between C–SAA and C–RO-D.
Extending these guarantees to bound worst-case regret or post-decision surprise under partial-feedback settings—where counterfactual costs are unobserved—remains an important direction for future work.
4.6 Application to the Omnichannel Multi-Courier CSOFP
The generic CSO framework applies directly to CSOFP via the following identifications:
$\mathcal{I}\;\mapsto\;\{\text{SKUs}\},$ $\mathcal{V}\;\mapsto\;\mathcal{K}\times\mathcal{L},$
$z_{iv}\;\mapsto\;z_{ik\ell},$ $\mathbf{s}_{v}\;\mapsto\;\mathbf{s}_{k\ell},$
$\tilde{\mathbf{c}}=(\tilde{c}_{v})_{v\in\mathcal{V}}\;\mapsto\;\tilde{\mathbf{d}}=(\tilde{d}_{k\ell})_{k\in\mathcal{K},\ell\in\mathcal{L}}.$ $\mathbb{P}_{v}\;\mapsto\;\mathbb{P}_{k\ell}.$
Under the conditional independence assumption,
$\mathbb{P}({\mathbf{d}}\mid\mathbf{s})=\prod_{k\in\mathcal{K}}\prod_{\ell\in\mathcal{L}}\mathbb{P}_{k\ell}(d_{k\ell}\mid\mathbf{s}_{k\ell}).$
Hence CSOFP (Problem 6) can be rewritten as
$\min_{\mathbf{z}\in\mathcal{Z}}\mathbb{E}_{\tilde{\mathbf{d}}\sim\Pi_{k,\ell}\mathbb{P}_{k\ell}(d_{k\ell}\mid\mathbf{s}_{k\ell})}[g(\mathbf{z},\tilde{\mathbf{d}})].$ (17)
To learn each marginal distribution $\mathbb{P}_{k\ell}(d_{k\ell}\mid\mathbf{s}_{k\ell})$, partition the historical dataset into
$\mathbb{D}_{k\ell}=\bigl{\{}(\mathbf{s}_{k\ell,m},\,d_{k\ell,m})\bigr{\}}_{m=1}^{M_{k\ell}},$
and train a per‑carrier model $\mathcal{M}_{k}$ on $\mathbb{D}_{k}=\bigcup_{\ell\in\mathcal{L}}\mathbb{D}_{k\ell}$. The full contextual distribution oracle is $\mathcal{M}(\mathbf{s})=\bigl{(}\mathcal{M}_{k}(\mathbf{s}_{k\ell})\bigr{)}_{k\in\mathcal{K},\,\ell\in\mathcal{L}}.$
This per‑carrier decomposition simplifies learning, generalizes to new locations, and outperforms a single global model according to preliminary experiments.
4.6.1 Applying C-SAA and C-RO
Computing the expectation in $\mathbb{E}_{\tilde{\mathbf{d}}\sim\mathcal{M}(\mathbf{s})}\bigl{[}g(\mathbf{z},\tilde{\mathbf{d}})\bigr{]}$ directly would require either enumerating all $C^{|\mathcal{K}|\times|\mathcal{L}|}$ deviation vectors or numerically integrating a continuous CDF—both computationally intractable. Instead, the two solution paradigms use the contextual distribution oracle $\mathcal{M}(\mathbf{s})$ as follows:
  • C-SAA replaces the expectation by a sample average: draw $N_{1}$ full‑vector scenarios $\{\tilde{\mathbf{d}}^{n}\}\sim\mathcal{M}(\mathbf{s})$ and solve
    $\min_{\mathbf{z}\in\mathcal{Z}}\;\frac{1}{N_{1}}\sum_{n=1}^{N_{1}}g\bigl{(}\mathbf{z},\tilde{\mathbf{d}}^{n}\bigr{)}.$
  • C-RO builds an uncertainty set from oracle outputs—either the discrete scenario set $\{\tilde{\mathbf{d}}^{n}\}_{n\in[N]}$ or a budgeted‑interval set from marginal quantiles—and then solves
    $\min_{\mathbf{z}\in\mathcal{Z}}\;\max_{\mathbf{d}\in\mathcal{U}(\mathbf{s})}g(\mathbf{z},\mathbf{d}).$
Both approaches avoid the exponential support growth while remaining agnostic to the internal structure of $\mathcal{M}$. A full C‑RO‑B reformulation is provided in the Appendix.
5 Learning the Contextual Distribution of Delivery Deviations for CSOFP
Having defined in Section 4 a generic contextual distribution oracle $\mathcal{M}(\mathbf{s})$, this section describes its concrete implementation for CSOFP. It begins by selecting features to construct the covariate vector. Two classes of machine learning methods are then proposed to estimate the contextual distribution of deviations: (1) Probabilistic multi-class classification, which leverages the discrete support of deviations; and (2) Tree-based quantile regression, which directly estimates the conditional CDF and avoids quantile crossing (Salari et al. 2022).
For simplicity, the carrier index $k$ is omitted in the notation, and the dataset for training the per-carrier model is written as $\mathbb{D}=\{(\mathbf{s}_{m},d_{m})\}_{m\in[M]}$.
5.1 Contextual Feature Selection
Carefully selecting contextual features is crucial for constructing informed covariates for delivery deviation predictions, as various stages in the order life cycle can contribute to deviations.
Candidate features fall into three categories:
  1. 1.
    Order-level: This includes customer location, SKU weight and dimensions, SKU type, quantities, release hour, release day, and planned lead time (i.e., the difference between the promised delivery date and the order release time). These factors can influence deviations through logistical complexity, handling requirements, and time-sensitive elements affecting the order’s processing and delivery schedule.
  2. 2.
@@ 5.1 Contextual Feature Selection
    Fulfillment center: This includes coordinates, location type, on-hand capacity, and on-hand inventory. The geographical location and type of fulfillment center can affect the efficiency of order processing, while capacity and inventory levels determine the speed and reliability of fulfilling orders.
@@
    Carrier-related: This includes service level, promised transit time, shipping charges, and carrier zone. These factors impact the carrier ability to meet delivery deadlines, with service level and transit times directly affecting deviations, while shipping charges and zones reflect logistical challenges and potential bottlenecks.
5.2 Probabilistic Multi-Class Classification
This discrete nature allows the contextual distribution learning problem to be framed as a multi-class classification (MC-CLF) problem. Under MC-CLF, each class represents a possible delivery deviation value in number of days, and the output class probabilities correspond directly to the probability mass function of the discrete distribution. For a new sample with covariates vector $\mathbf{s}$, let the conditional probability of class $c$ be $p_{c}(\mathbf{s})=\mathbb{P}(d=\xi_{c}\mid\mathbf{s})$, with the corresponding predicted probability from MC-CLF denoted as $\hat{p}_{c}(\mathbf{s})$.
Multinomial logistic regression (MLR) and classification tree-based models are considered to solve MC-CLF. Specifically, MLR is a classic classifier that aims to estimate the class probabilities by maximizing the log-likelihood function, the negative of the log loss function. The log loss is defined as $-\sum_{m=1}^{M}\sum_{c=1}^{C}\mathbf{1}\{d_{m}=\xi_{c}\}\log(\hat{p}_{c}(\mathbf{s}_{m})).$ On the other hand, classification trees use a nonparametric approach that recursively partitions the data into subsets based on feature values in a top-down manner, resulting in a tree-like model of decisions.
Let $\mathcal{R}_{1},\ldots,\mathcal{R}_{R}$ be the regions partitioned by a classification tree. Classification tree estimates
@@ 5.2 Probabilistic Multi-Class Classification
The performance of the MC-CLF models are evaluated by log loss and Brier score. The latter is defined as the average squared error between predicted probabilities and actual outcomes $\sum_{m=1}^{M}\sum_{c=1}^{C}(\mathbf{1}\{d_{m}=\xi_{c}\}\hat{p}_{c}(\mathbf{s}_{m}))^{2}$. For both metrics, lower values indicate better probabilistic prediction performance.
5.2.1 Probability Calibration
Raw class‐probability estimates $\hat{p}_{c}(\mathbf{s})$ from classifiers often exhibit mis‐calibration: a predicted probability $\alpha$ does not correspond to an empirical frequency $\alpha$ (Niculescu-Mizil and Caruana 2005). Good calibration means that, over all samples with $\hat{p}_{c}(\mathbf{s})=\alpha$, the true fraction belonging to class $c$ is approximately $\alpha$. Formally, let $f(\alpha)=\mathbb{P}\bigl{(}d=\xi_{c}\mid\hat{p}_{c}(\mathbf{s})=\alpha\bigr{)}$ be the calibration function. Perfect calibration means $f(\alpha)=\alpha$ for all $\alpha\in[0,1]$.
Isotonic regression (Zadrozny and Elkan 2002) learns a non‐decreasing mapping $g:[0,1]\to[0,1]$ that best aligns raw scores to observed frequencies, preserving the rank order of $\hat{p}_{c}$. Given a calibration set $\mathbb{D}_{cal}=\{(\hat{p}_{c}(\mathbf{s}_{m}),\,y_{c,m})\}_{m=1}^{M_{cal}}$ with $y_{c,m}=\mathbf{1}\{d_{m}=\xi_{c}\}$, isotonic regression solves
$\min_{g}\;\sum_{m=1}^{M_{cal}}\bigl{(}y_{c,m}-g(\hat{p}_{c}(\mathbf{s}_{m}))\bigr{)}^{2}\quad\text{s.t.}\quad g(\alpha)\leq g(\alpha^{\prime})\;\text{whenever}\;\alpha\leq\alpha^{\prime}.$
The calibrated probabilities are $\tilde{p}_{c}(\mathbf{s})=g(\hat{p}_{c}(\mathbf{s}))$. By enforcing monotonicity, isotonic calibration ensures that higher raw scores remain higher after mapping, and empirical frequencies within each score bin match the calibrated values.
In this paper, a $k$-fold cross‐validation procedure is employed: for each fold $i$, the classifier is trained on its training split, raw scores are obtained on the held‐out calibration split, and an isotonic regression model $g^{(i)}$ is fit to those scores. On unseen test instance with covariates $\mathbf{s}$, each fold’s classifier and corresponding $g^{(i)}$ yield calibrated probabilities $\tilde{p}_{c}^{(i)}(\mathbf{s})$, which are then averaged across all $k$ folds to produce the final estimate $\tilde{p}_{c}(\mathbf{s})=\frac{1}{k}\sum_{i=1}^{k}\tilde{p}_{c}^{(i)}(\mathbf{s})\,.$
5.3 Tree-Based Quantile Regression
Unlike multi‑class classification methods that treat each deviation value as an unordered category, quantile regression (QR) exploits the inherent ordering of deviations by framing the task as ordinal regression and directly estimating the conditional cumulative distribution function (CDF) $F(u\mid\mathbf{s})=\mathbb{P}(d\leq u\mid\mathbf{s})$.
Standard QR models minimizes, for each quantile level $q$, the pinball loss
$\rho_{q}(d,\hat{d})\;=\;q\,\max(d-\hat{d},0)\;+\;(1-q)\,\max(\hat{d}-d,0),$ (19)
where $\hat{d}$ denotes the model’s predicted $q$‑th quantile. Collecting these quantile estimates across $q$ yields a stepwise approximation of the CDF. However, estimating each quantile independently can violate the ordering $\hat{d}_{(q_{1})}\leq\hat{d}_{(q_{2})}$ for $q_{1}<q_{2}$, resulting in a non‑monotonic—and thus invalid—CDF.
Tree‐based QR methods—namely regression trees extended to CDF estimation, and their ensemble version, quantile random forests (QRF) (Meinshausen and Ridgeway 2006)—avoid this “quantile crossing” problem by extracting quantiles from the empirical distribution of training sample in each leaf region.